In a right-angled triangle, one of the sides adjacent to the right angle is 4cm longer than the other side
Find the length of ALL three sides if the area of the triangle is 96cm(squared)

Answer by mananth(16946)

1/2 * base * height = 96
base * height = 192
Let one side be x
the other be (x+4)
....
x(x+4)=192
x^2+4x=192
x^2+4x-192=0
x^2+16x-12x-192=0
x(x+16)-12(x+16)=0
(x+16)(x-12)=0
x= 12 cm one side
other side = 16 cm
12^2+16^2=HYP^2
144+256=HYP^2
400=HYP^2
HYP = 20 cm